Session: May 2001 (session 2)
Wavelength: 18. cm
To give feedback on this experiment, click
\(Other\) (A. Cattani at Tue May 29 05:26:28 METDST 2001) Success
Torun (Kazimierz Borkowski at Tue May 29 09:37:22 METDST 2001) Minor failures
No problems till 0:14 UTC (on day 148), when power failure prevented
further observations (so the 3rd tape has only ~1 hour of data).
Westerbork (Tony Foley at Tue May 29 11:49:10 METDST 2001) Minor failures
Westerbork (Tony Foley at Tue May 29 11:49:51 METDST 2001) Minor failures
Some EMI. Otherwise it looks OK.
Effelsberg (A. Kraus at Tue May 29 13:13:11 METDST 2001) Minor failures
Some scans were lost - not enough slewing time. HPE on forward tracks.
Noto (P. Leto at Wed May 30 13:15:21 METDST 2001) Success
No known problem
Jodrell_Bank (Alastair Gunn at Wed May 30 17:56:23 METDST 2001) Success
No known problems.
Onsala (A. Polatidis at Thu May 31 11:45:40 METDST 2001) Success